umerlodhi (umerlodhi):

Which method is the most efficent method to use to solve x^2-11=0? factoring isolating the x^2 term and finding the square root of both sides using the quadratic formula all three methods would be efficent

umerlodhi (umerlodhi):

@ganeshie8 @TheSmartOne

umerlodhi (umerlodhi):

@jabez177 @sleepyjess

OpenStudy (sleepyjess):

What do you think? Which option will be quickest and easiest?

umerlodhi (umerlodhi):

no wait i think its B

OpenStudy (sleepyjess):

That's not what I was thinking, the quadratic formula takes some time to work out

OpenStudy (sleepyjess):

Yep :)

umerlodhi (umerlodhi):

thanks so much!!

OpenStudy (sleepyjess):

No problem ^_^

TheSmartOne (thesmartone):

IMO, I think the answer is D Factoring it using: a^2 - b^2 = (a + b)(a - b) Using the quadratic formula or isolating x^2 x^2 = 11 x = plus or minus sqrt(11)

OpenStudy (sleepyjess):

@TheSmartOne , this equation is impossible to solve by factoring though

OpenStudy (sleepyjess):

It also says the MOST efficient, I wouldn't consider the quadratic formula to be "efficient" for this equation

TheSmartOne (thesmartone):

I guess that's what they want you to say :P factoring for this question is efficient if they teach kids a^2 - b^2 = (a + b)(a - b)

OpenStudy (sleepyjess):

but 11 is prime, it can't be factored

TheSmartOne (thesmartone):

\(\Large (\sqrt{11})^2 = 11\)

OpenStudy (sleepyjess):

using \(\sqrt{11}\) isn't exactly ideal for factoring....

TheSmartOne (thesmartone):

The correct answer is actually D I helped someone with it and they got D as the correct answer

umerlodhi (umerlodhi):

oh snap i submitted it yesterday and got 4/5 idk which one was wrong and now the quiz is locked so i cant check